Majer Corporation makes a product with the following standard costs: 






Standard Quantity or HoursStandard Price or  RateStandard Cost Per Unit
Direct materials6.1 ounces$3.00 per ounce$18.30
Direct labor0.4 hours$16.00 per hour$6.40
Variable overhead0.4 hours$3.00 per hour$1.20


The company reported the following results concerning this product in February 

Originally budgeted output5,000 units
Actual output6,100 units
Raw materials used in production33,500 ounces
Actual direct labor-hours2,070 hours
Purchases of raw materials35,900 ounces
Actual price of raw materials$57.10 per ounce
Actual direct labor rate$47.60 per hour
Actual variable overhead rate$5.70 per hour


The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased. 

The variable overhead efficiency variance for February is:
Answer #1

Variable overhead efficiency variance

= (SLH - ALH) *SVOHR

= (6100*0.4 - 2070) * 3

= (2440 - 2070) * 3

= 1110 Favourable

Comment if you face any issues